Exercise 16-15

Sweet Inc. uses a calendar year for financial reporting. The company is authorized to issue 8,420,000 shares of $12 par common stock. At no time has Sweet issued any potentially dilutive securities. Listed below is a summary of Sweet’s common stock activities.

1. Number of common shares issued and outstanding at December 31, 2018

2,010,000

2. Shares issued as a result of a 12% stock dividend on September 30, 2019

241,200

3. Shares issued for cash on March 31, 2020

2,060,000

Number of common shares issued and outstanding at December 31, 2020

4,311,200

4. A 2-for-1 stock split of Sweet’s common stock took place on March 31, 2021

In 2019, Flounder Enterprises issued, at par, 60 $1,000, 8% bonds, each convertible into 100 shares of common stock. Flounder had revenues of $19,500 and expenses other than interest and taxes of $9,500 for 2020. (Assume that the tax rate is 20%.) Throughout 2020, 1,600 shares of common stock were outstanding; none of the bonds was converted or redeemed.

(a) Compute diluted earnings per share for 2020. (Round answer to 2 decimal places, e.g. $2.55.)

Earnings per share $

Entry field with incorrect answer now contains modified data


(b) Assume the same facts as those assumed for part (a), except that the 60 bonds were issued on September 1, 2020 (rather than in 2019), and none have been converted or redeemed. Compute diluted earnings per share for 2020. (Round answer to 2 decimal places, e.g. $2.55.)

Earnings per share $

Entry field with incorrect answer now contains modified data


(c) Assume the same facts as assumed for part (a), except that 20 of the 60 bonds were actually converted on July 1, 2020. Compute diluted earnings per share for 2020.
Answer #1

Please find below table useful to compute desired results: -

в Particulars given 60 1000 =8/100 100 shares 19500 9500 =20/100 1600 -B7 1 2 Bonds issue 3 Par Value of Bond 4 Interest RateА 28 =B6 29 (C) Revenues 30 Expenses 31 Other than interest 32 Bond Interest 33 Bond Interest 34 Income before Income taxes 3

End results would be as follows: -

BC 1 shares А Particulars given 2 Bonds issue 3 Par Value of Bond 4 Interest Rate 5 Convertible into 6 Revenues 7 Expenses 8B C | 19,500 28 29 (c) Revenues 30 Expenses 31 Other than interest 32 Bond Interest 33 Bond Interest 34 Income before Income